OpenStudy (anonymous):

Match each proportion with its solution. Match Term Definition five over x equals fifteen over six A) eighteen twelve over sixteen equals x over forty B) twelve four over twelve equals six over x C) thirty x over eighteen equals eight over twelve D) 2

OpenStudy (texaschic101):

5/x = 15/6 cross multiply 15x = 30 x = 2 (D) 12/16 = x/40 cross multiply 16x = 480 x = 30 (C) 4/12 = 6/x cross multiply 4x = 72 x = 18 (A) x/18 = 8/12 cross multiply 12x = 144 x = 12 (B)

OpenStudy (anonymous):

please help me look at the attachment

OpenStudy (texaschic101):

do you have any questions about anything that I did ?

OpenStudy (anonymous):

how do you set it up and work it out?

OpenStudy (texaschic101):